# Do graph neural network states contain graph properties?

### v1

@@ 5 Methodology
For each of the three datasets, we use a similar network architecture consisting of a number of GNN (GCN, GIN, or GAT) layers, followed by a pooling operation (mean- (Kipf & Welling, 2017), sum-(Xu et al., 2019), or max-pooling (Hamilton et al., 2017)), and then a number of dense layers. We optimize the hyperparameters to obtain good models for the binary classification task.
@@
For the Grid-House dataset the hyperparameter information can be found in table 4. On this dataset we also compared different regularisation methods. The explicit $L_{2}$ regularisation encourages the network to keep the weights small, and we expect that this would make the embeddings less sensitive to fluctuations in the input data and smoother. the latter would make them more linearly separable for our probing methods. Dropout randomly disables a fraction of the neurons during each training iteration which forces the network to learn redundant representations, as any neuron could be dropped out. These redundant representations might make it more difficult to linearly separate the graph embeddings. We ran each model 20 times and took the one with the best accuracy.
@@
For the ClinTox Molecular dataset, we ranged the number of layers from 4 to 6 and hidden dimensions from 64 to 256. The final model architectures were selected based on optimal performance on the ClinTox dataset.
For fMRI FC connectomes the hyperparamter search space is described in table 15
5.1 Probing strategy
@@
Probing is performed on the train and test sets, where train features $\{f_{\text{train}}^{(i)}\}$ and graph properties $\{z_{\text{train}}^{(i)}\}$ are paired for each graph (equally for the test set). Let’s define at least one example for the GCN model. Let $\mathcal{G}^{i}=(A^{i},X^{i})$ denote the $i$-th graph, where $A^{i}$ is the adjacency matrix and $X^{i}$ is the node feature matrix as previously defined. The GCN layers iteratively update the node features $H^{(l)}$ through graph convolutions defined previously as $H^{(l+1)}=\sigma(\hat{A}H^{(l)}W^{(l)})$, where $\hat{A}$ is the normalized adjacency matrix, $W^{(l)}$ are the trainable weights, and $\sigma$ is a non-linear activation function (ReLU). The node embeddings $H^{(l)}$ at each layer $l$ capture both local and global structural information by aggregating features from neighboring nodes. The final node embeddings $H^{(4)}$ are pooled using global max pooling to generate a graph-level embedding $H_{\text{global}}$, which is passed through three fully connected layers to produce the final prediction $\hat{y}$. We deinfe these post pooling operations as $H_{\text{global}}^{(5)}=\sigma(W_{1}H_{\text{global}}),H_{\text{global}}^{(6)}=\sigma(W_{2}z_{1}),\hat{y}=\text{Softmax}(W_{3}z_{2})$. For probing purposes, we use $H^{(l)}$ at different layers to evaluate node-level properties, while $H_{\text{global}},H_{\text{global}}^{(5)},H_{\text{global}}^{(6)}$, and $\hat{y}$ are used to assess graph-level properties.
@@
To address this, we first sort the embeddings in descending order based on their norms before concatenating, which introduces permutation invariance. Sorting in this way ensures that any padding zeros align at the end of the sequence, enabling learnable representations for graphs with varying node counts. While sorting for permutation invariance is not widely discussed in the literature, it provides a practical solution by using the embeddings’ properties to enforce consistent ordering across graphs.
@@ 6.1 Grid-House dataset
The models performed as anticipated thanks to their high expressiveness and the linearly separable nature of the classification problem as we can see in table 5. The probing results on the Grid-House dataset demonstrate that the number of squares consistently yields the highest $R^{2}$ scores across all models in the global graph embeddings (after pooling aggregation has been applied). This aligns with our initial hypothesis.
In general, higher-layer embeddings filter out many other graph properties as they are less relevant for making the classification problem linearly separable. The GNN’s final layer focuses on the number of squares, effectively partitioning the graphs into two classes: those with #$\#\text{squares}<5$ (indicating either the grid or house alone) and those with #$\#\text{squares}=5$ (indicating the presence of both substructures). This reduction in feature space through the layers aligns with the model’s goal of optimizing the decision boundary for binary classification, where the number of squares becomes a clear and dominant factor for separability. Further confirming expectations, density and average path length are also prominent as the presence of both a house and a grid does slightly increase the average density and path length of graphs. These findings confirm the correspondence between graph embeddings clustering and property hyperplane separation as shown in fig. 5.
@@
We further observe that, for both the GCN and GIN models, the application of $L_{2}$ regularization yields the expected behavior. The last layer of the GCN in fig. 7 shows a stronger dominance of the number of squares feature when $L_{2}$ regularization is applied compared to when it is not. Similarly, in the GIN, both number of triangles and density become less detectable relative to the number of squares, by the probing classifier in the final layers under $L_{2}$ regularization, consistent with the anticipated effects on the feature representation.
We observe results consistent with our expectations for the models with dropout. The key property is less dominant, and multiple properties are represented in the final layers. Notably, in fig. 8 the last layer, the separability gap between the #square and the other properties is reduced, indicating a more distributed representation of features when dropout is applied.
@@
The GAT model stands out by capturing not only squares ($R^{2}=0.88$), but also performing well on other properties like triangles, cliques, and density. GAT assigns a weight to each neighbouring node based on a learned function of the node features, aggregating the neighbours’ information in a weighted manner. This feature-dependent mechanism introduces flexibility but also makes GAT’s performance contingent on the quality and richness of the node features. It seems that GAT’s broader capability compared to the GCN comes at the cost of focus, as GAT tends to incorporate multiple features, which may dilute its ability to pinpoint the most crucial property (in this case, the number of squares) for the classification task. This over-reliance on feature aggregation can lead to inefficiencies when simpler, more targeted properties suffice, as seen with GIN. These results also make sense with regard to the best score obtained by the GIN architecture as seen in table 5.
@@ 6.2 ClinTox Molecular
As expected GIN outperformed the other models. Based on scores and good properties (inferential mechanism with better expressivity) we focus on the GIN results for the results on other datasets. Detailed results can be found in the appendix, table 8. When looking into the linear probing performance in table 1, we find that the highest scores are consistently yielded by the average degree, the spectral radius, the algebraic connectivity and the density, in that order.
@@
The average degree of atoms in a molecule provides a straightforward interpretation, as atoms with higher valencies are generally less stable and less biologically compatible. For instance, hydrogen with a valency of 1 and oxygen with a valency of 2 are more compatible with carbon-based molecules, whereas sulfur, with a valency of 6, is less favorable for biological systems  (Komarnisky et al., 2003). Therefore, the average degree serves as a useful indicator of molecular toxicity. Additionally, the spectral radius, often associated with molecular stability and reactivity, is another valuable graph property. Molecules with a lower spectral radius tend to be more stable, while those with a higher spectral radius may exhibit localized electron densities, increasing their reactivity. Using this property to predict molecular toxicity is a logical approach. To the best of our knowledge, there is no comprehensive analysis exploring the role of spectral radius in the emergence of molecular toxicity, highlighting an opportunity for future research.
@@
In the detailed results, in the appendix table 14, GIN outperforms the other architectures in both parts of the dataset (reproducing the observation by Zheng et al. (2023). Again, the strength of GIN lies in its injective aggregation mechanism. The probing results on the ASD dataset reveal that the number of triangles consistently achieves high $R^{2}$ scores across all models, with particularly strong performance in GIN models. This property is followed by the spectral radius and the density.
As further detailed in the appendix tables 16, 18, 19 and 17 the number of edges is particularly well encoded in the representation of the GAT. This is a consequence of its reliance on feature-dependent weights and structure-free normalisation, which limit its ability to capture specific structural properties that do not directly depend on edges. The GCN results are broadly comparable to those of the GIN, though they tend to be less precise and selective.
For the MDD results we also focus the GIN model. Detailed results are in the appendix tables 20, 23, 24 and 22. The probing results MDD reveal that the number of triangles still consistently achieves high $R^{2}$ scores across all models while being less of a distinctive feature than in ASD. This time, the spectral radius is dominated by the density of the graph. In general, the embeddings from the 7th layer of our GIN architecture exhibit higher $R^{2}$ scores for relevant graph properties, suggesting improved separability in the embedding space for MDD classification compared to ASD. This indicates that the learned representations at this depth capture more discriminative structural features, facilitating more effective class separation between MDD and healthy controls.
@@ 7.1 Expectations
For Grid-House we hypothesized that the GNN would benefit from leveraging both the local clustering coefficient and eigenvector centrality as node-level features. The first one would help characterize a house, the second a grid. However, neither feature alone is sufficient to render the problem linearly separable. We therefore expect either a combination of features or a single global property (e.g. number of squares) to be leveraged. If the tensor embeddings produced by the GNN can be used to predict these properties, this would indicate that the GNN is utilising them in solving the classification task.
For the ClinTox Molecular dataset, based on the literature Kengkanna & Ohue (2024); Chen et al. (2021); Jiang et al. (2021) some few properties have been found to be link with toxicity such as the node degree (i.e. the valency), subgraph patterns (functional groups, chemical fragments), and the overall graph connectivity.
Based on existing literature on functional connectivity (FC) network properties in ASD and MDD, we hypothesized that specific properties will be critical in classifying brain networks for the fMRI FC connectomes dataset. For ASD, we expect betweenness centrality to play a significant role at the node level, reflecting local overconnectivity. At the graph level, we anticipate that clustering coefficient, characteristic path length, and small-worldness will be essential in capturing the local and global network disruptions seen in ASD, particularly the imbalance between local overconnectivity and long-range underconnectivity. For MDD, we hypothesise that increased clustering coefficients, modularity, number of triangles and number of squares will be key features for classification, as they could indicate of heightened local interconnectedness and disrupted global integration.
@@ 7.2 Findings
We first demonstrate the feasibility of our probing method through the Grid-House dataset. This acts as a proof of concept on probing classifiers plumbing the representations learned by the GNN. We made sure to choose a classification task which requires learning global structural properties of the graph, such as motifs like squares and triangles, and a long range dependencies between those. In line with our expectations, the results show significant dominance of number of squares in the post pooling layers of every model, while still highlighting the superiority of the GIN model in leveraging superior representations when trained to classify graphs with complex motifs like Grid and House. This shows the expressivity of the models and their ability to reduce the complexity of a graph related problem to known graph properties, making the problem linearly separable in the space of their embeddings. These results are consistently higher than those obtained from probing the models with randomized labeling, highlighting the relevance of the initial findings.
@@
Using the ClinTox Molecular dataset to assess molecular toxicity, we explored how key graph properties, such as the average degree and spectral radius, are utilized by our GIN architecture. The average degree, closely linked to atomic valency, reflects a molecule’s potential for interactions. The spectral radius offers a complementary hypothesis, suggesting that the overall structural stability of a molecule, independent of specific atomic features, may also be a key factor in toxicity prediction. These results suggest that despite the limitations of our approach, it still holds potential for assisting research in complex systems fields, such as neuroscience or social sciences, where emergent phenomena play a crucial role in understanding system dynamics.
@@
Given the previous positive results, we explored a real life applications with the fMRI FC connectomes dataset. Here, the results provide new insights that extended beyond our initial hypothesis. While we expected betweenness centrality, clustering coefficient, and characteristic path length and Small-worldness to be the most relevant for distinguishing ASD from healthy individuals, the prominence of the number of triangles highlighted the importance of local structural motifs. This makes sense in the context of functional connectivity, where local overconnectivity in specific brain regions, such as sensory and association cortices, has been observed in individuals with ASD. The strong role of triangle motifs may reflect the tight, redundant local connections that characterize these regions, supporting the hypothesis that local overconnectivity is a key factor in ASD. The spectral radius and density and graph energy being particularly significant is also logical, as these properties are closely related to the overall connectivity strength and the compactness of connections within subnetworks. The presence of SW is still quite significant in the post-pooling layers which is interesting. In ASD, where global integration is often reduced and local connectivity heightened, these metrics may provide an important reflection of the imbalance between short-range and long-range communication pathways in the brain.
The outcomes for both the ASD and the MDD datasets showed promising results that should be discussed more deeply with neuroscientists. The results mainly suggest the importance of graph substructures or spectral and small-world properties over more basic graph properties to explain how graph neural networks predict these neurological disorders in the FC matrices of patients. The presence of algebraic connectivity in the last two layers of the GIN MDD probing setting while being completely absent from the GIN ASD probing one makes it a property of choice to examine for understanding better how this kind of neurological disorders affect the brain’s connectivity. This is a result which, to the best of my knowledge, has never been investigated before.
7.3 Comparison between datasets

### v2

@@ 4 Methodology
For each of the three datasets, we use a similar network architecture consisting of a number of GNN (GCN, GIN, or GAT) layers, followed by a pooling operation (mean- (Kipf & Welling, 2017), sum-(Xu et al., 2019), or max-pooling (Hamilton et al., 2017)), and then a number of dense layers. We optimized the hyperparameters to obtain the best models possible. For the Grid-House dataset the complete hyperparameter information can be found in table 4. We ran each model 20 times and took the one with the best accuracy. As expected, there is a correlation of 0.992 between task accuracy and the maximum probing accuracy as seen in fig. 6.
@@
We compared different regularization methods. The explicit $L_{2}$ regularization encourages the network to keep the weights small, and we expect that this would make the embeddings less sensitive to fluctuations in the input data which would translate in later layers being more selective to graph properties. Dropout randomly disables a fraction of the neurons during each training iteration which forces the network to learn redundant representations, as any neuron could be dropped out. These redundant representations might make it more difficult to linearly separate the graph embeddings. We expect later layers to distinguish less between graph properties. We plot only post-pooling layers for the sake of clarity.
@@
For the ClinTox Molecular dataset, we ranged the number of layers from 4 to 6 and hidden dimensions from 64 to 256. The final model architectures were selected based on optimal performance on the ClinTox dataset. For fMRI FC connectomes the hyperparamter search space is described in table 15.
Figure 3: Visualization of test set graph in ASD dataset with node sizes based on (from left to right): (a) degree, (b) betweenness centrality, (c) eigenvector centrality, and (d) clustering coefficient.
@@
Probing is performed on the train and test sets, where train features $\{f_{\text{train}}^{(i)}\}$ and graph properties $\{z_{\text{train}}^{(i)}\}$ are paired for each graph (equally for the test set). Let’s define at least one example for the GCN model. Let $\mathcal{G}^{i}=(A^{i},X^{i})$ denote the $i$-th graph, where $A^{i}$ is the adjacency matrix and $X^{i}$ is the node feature matrix as previously defined. The GCN layers iteratively update the node features $H^{(l)}$ through graph convolutions defined previously as $H^{(l+1)}=\sigma(\hat{A}H^{(l)}W^{(l)})$, where $\hat{A}$ is the normalized adjacency matrix, $W^{(l)}$ are the trainable weights, and $\sigma$ is a non-linear activation function (ReLU). The node embeddings $H^{(l)}$ at each layer $l$ capture both local and global structural information by aggregating features from neighboring nodes. The final node embeddings $H^{(4)}$ are pooled using global max pooling to generate a graph-level embedding $H_{\text{global}}$, which is passed through three fully connected layers to produce the final prediction $\hat{y}$. We define these post pooling operations as $H_{\text{global}}^{(5)}=\sigma(W_{1}H_{\text{global}}),H_{\text{global}}^{(6)}=\sigma(W_{2}z_{1}),\hat{y}=\text{softmax}(W_{3}z_{2})$. For probing purposes, we use $H^{(l)}$ at different layers to evaluate node-level properties, while $H_{\text{global}},H_{\text{global}}^{(5)},H_{\text{global}}^{(6)}$, and $\hat{y}$ are used to assess graph-level properties.
@@
To address this, we compare two methods. One where we mean-pool the node embeddings before probing them like in fig. 12. But we might miss relevant basic properties. At the same time, we make sure that our probe is not extrapolating more structural information that what is encoded in juxtaposed node embeddings. A probing classifier, by operating on the set of embeddings simultaneously, could learn to infer relationships and patterns between nodes, leveraging the contextual relationships implicit in their embeddings. The other one where we first sort the embeddings in descending order based on their norms before concatenating fig. 4. This ordering depends only on the inherent properties of the embeddings themselves, not on their original ordering in the graph. As such, it inherently respects permutation invariance because reordering the nodes does not affect their norms or the resulting sorted order. Conveniently, sorting in this way ensures that any padding zeros align at the end of the sequence, enabling learnable representations for graphs with varying node counts. While sorting for permutation invariance is not widely discussed in the literature, it provides a practical solution by using the embeddings’ properties to enforce consistent ordering across graphs.
@@ 5.1 Grid-House dataset
The models achieved high performance scores, ranked according to their expressivity: GCN scored 0.90, GAT 0.97, and GIN 1.00 (table 5). The probing results demonstrate that the number of squares consistently yields the highest $R^{2}$ scores across all models in the global graph embeddings, both with mean-pooled GNN layer and without. This confirm our initial hypothesis where the number of square is the property of interest to perform this classification, as show the correlation between the highest R2 scores (# squares) and the performance of models fig. 6.
@@
When comparing the mean-pooling and norm-sorting methods, the key observation is that mean-pooling renders the probe unable to predict basic properties such as the number of nodes. It significantly reduces the probe’s ability to predict clustering and centrality-related properties and slightly diminishes its performance in predicting higher-order properties, such as the # square.
In general, the spectral and spatial perspectives are confirmed as deeper layers further smooth the signal, capturing global properties but potentially losing high-frequency details. Analogically, they extend their receptive field, it’s well visible with the message passing GCN fig. 8, incorporating more distant nodes into each node’s representation as k-hop neighborhood are taken into account. As seen in the T-SNE visualizations figs. 8, 12 and 17, the final layers tend to focus on the # squares, effectively partitioning the graphs into two classes: those with #$\#\text{squares}<5$ (indicating either the grid or house alone) and those with #$\#\text{squares}=5$ (indicating the presence of both substructures). This reduction in feature space through the layers aligns with the model’s goal of optimizing the decision boundary for binary classification, where the number of squares becomes a clear and dominant factor for separability. Interestingly enough, the GIN model has its layer MLP1 partitioning the graphs label 0 between those with a house and those with a grid using # of triangle significantly. The Gat model present a form of confusion between classes at the layer Global before clearly sharping its representation and classification abilities. The other present properties, density and average path length are also prominent and make sense as the presence of both a house and a grid does slightly increase the average density and path length of graphs. These findings confirm the correspondence between graph embeddings clustering and property hyperplane separation.
We further observe that, for both the GCN and GIN models, the application of $L_{2}$ regularization yields the expected behavior. The last layer of the GCN in fig. 10 shows a stronger dominance of the number of squares feature when $L_{2}$ regularization is applied compared to when it is not. Similarly, in the GIN fig. 14, both # triangles and density become less detectable relative to the number of squares, by the probing classifier in the final layers under $L_{2}$ regularization, consistent with the anticipated effects on the feature representation. Dropout leads to the representation of multiple properties in the final layers. Notably, in fig. 11, the last layer exhibits a reduced separability gap between the # square and other properties, indicating a more distributed feature representation when dropout is applied.
We can use the number of square R² score of models to compare their expressivity. For the GCN fig. 8, the slow increase of the R² score correspond to the message-passing theory as we said. The use of the density as a side property make sense but also indicate the lack of optimization of this model, which we can attest by looking at the T-SNE of the last layer where some graphs are misgrouped. The regularization seems to correct this effect partialy fig. 10.
The GAT and GIN models shows a strong dominance of # of square from the beginning. Either with an Attention-Weighted Aggregation or a MLP injective aggregation, we don’t expect such score from the first layer. The receptive field is limited to 1-hop nodes. This could be explained by a form of signal short-cut where local features like the degree of neighbors is already enough to expect a four-cycle pattern. Similarly, the closeness centrality of a node or its clustering coefficient is barely encoded in the first layers table 7. The effectivity of both the GAT and the GIN models would then catch these short-cuts very effectively.
@@ 5.2 ClinTox Molecular
As expected, the GIN model outperform the other models with a test accuracy of 0.93 table 8. When looking at the corresponding linear probing performance in table 1, we find that the highest scores are consistently yielded by the average degree, the spectral radius, the algebraic connectivity and the density, in that order.
@@
These findings validate our methodology on already known domain knowledge. Indeed, the average degree of atoms in a molecule provides a straightforward interpretation, as atoms with higher valencies are generally less stable and less biologically compatible. For instance, hydrogen with a valency of 1 and oxygen with a valency of 2 are more compatible with carbon-based molecules, whereas sulfur, with a valency of 6, is less favorable for biological systems  (Komarnisky et al., 2003). Therefore, the average degree serves as a useful indicator of molecular toxicity. Additionally, the spectral radius, often associated with molecular stability and reactivity, is another valuable graph property. Molecules with a lower spectral radius tend to be more stable, while those with a higher spectral radius may exhibit localized electron densities, increasing their reactivity. Using this property to predict molecular toxicity is a logical approach.
@@
Here, our primary focus is to uncover properties under exploited by the field. GIN outperforms other architectures table 14, reproducing the observation by Zheng et al. (2023). It makes sense to study exclusively the last layers of the GIN model. The probing results on the ASD dataset reveal that the number of triangles consistently achieves high $R^{2}$ scores across all models, with particularly strong performance in GIN models. This property is followed by the spectral radius and the density.
For the MDD results, detailed results tables 20, 23, 24 and 22 reveal that the number of triangles still consistently achieves high $R^{2}$ scores across all models while being less of a distinctive feature than in ASD. This time, the spectral radius is dominated by the density of the graph. In general, the embeddings from the 7th layer of our GIN architecture exhibit higher $R^{2}$ scores for relevant graph properties, suggesting improved separability in the embedding space for MDD classification compared to ASD. This indicates that the learned representations at this depth capture more discriminative structural features, facilitating more effective class separation between MDD and healthy controls.
@@ 6.1 Expectations
With Grid-House we hypothesized that the GNNs would benefit from leveraging the # of square to render the problem linearly separable. Based on their mathematical restrictions, we hypothesized that the GIN would perform better than the GAT and the GCN. Regularization methods should either refine the representations, as seen with $L_{2}$ regularization, or distribute them more broadly, as achieved with dropout.
Based on the message-passing paradigm, we anticipated a clear absence of the # of square in the first layer. Additionally, we expected that the mean-pooling and norm-sorting methods would not significantly alter how representations are probed, except for basic properties like the number of nodes, which are easily interpretable from the tensor of node vectors but not from an aggregated representation.
For the ClinTox Molecular dataset, based on the literature (Kengkanna & Ohue, 2024; Chen et al., 2021; Jiang et al., 2021) some few properties have been found to be link with toxicity such as the node degree (i.e. the valency), subgraph patterns (functional groups, chemical fragments), and the overall graph connectivity.
Based on existing literature on functional connectivity (FC) network properties in ASD and MDD appendix G, we hypothesized that specific properties would be critical in classifying brain networks for the fMRI FC connectomes dataset. For ASD, we expect betweenness centrality to play a significant role at the node level, reflecting local overconnectivity. At the graph level, we anticipate that clustering coefficient, characteristic path length, and small-worldness will be essential in capturing the local and global network disruptions seen in ASD, particularly the imbalance between local overconnectivity and long-range underconnectivity. For MDD, we hypothesise that increased clustering coefficients, modularity, number of triangles and number of squares will be key features for classification, as they could indicate of heightened local interconnectedness and disrupted global integration.
@@ 6.2 Findings
We first demonstrate the feasibility of our probing method through the Grid-House dataset. In line with our expectations, The number of squares metric dominated across all layers and models, with GIN showing enhanced expressivity. Secondary properties like avg_path_lenght figs. 7 and 8 showed early significance but gradually diminished through the layers, demonstrating how GNNs act as low-pass filters on graph signals, the receptive field select the most discriminative properties. These findings align with the labels, as the density or the average path length is increase when both a grid and a house are added to a graph. Mean-pooling and norm-sorting methods exhibit different behaviors that can be explained by the type of information the classifier access : aggregated graph data or detailed node embeddings. The probe might also infer relationships and patterns that are implicitly encoded in individual node embeddings but become explicit when considered collectively.
The analogy with computer vision lead to the question of supervision signal as we would expect unsupervised model to capture structural information more effectively than supervised ones (Kingma & Welling, 2022). Unsupervised learning algorithms can discover hidden patterns and structures in data without explicit instructions, potentially uncovering relationships that may not be immediately obvious to humans. Following this logic, this ability to find latent structures could be advantageous in capturing complex structural information in graph data and lead to similar breakthought than discovering new particle physics phenomena using clustering and dimensionality reduction on collider data or identifying subtypes of cancer with clustering and PCA in genomic studies.
@@
Using the ClinTox Molecular dataset to assess molecular toxicity, we explored how key graph properties, such as the average degree and spectral radius, are utilized by our GIN architecture. The average degree, closely linked to atomic valency, reflects a molecule’s potential for interactions. The spectral radius offers a complementary hypothesis, suggesting that the overall structural stability of a molecule, independent of specific atomic features, may also be a key factor in toxicity prediction.
@@
With the fMRI FC connectomes dataset, the results provided new insights that extended beyond our initial hypothesis. The prominence of the number of triangles highlighted the importance of local structural motifs. This makes sense in the context of functional connectivity, where local overconnectivity in specific brain regions, such as sensory and association cortices, has been observed in individuals with ASD. The strong role of triangle motifs may reflect the tight, redundant local connections that characterize these regions, supporting the hypothesis that local overconnectivity is a key factor in ASD. The spectral radius and density and graph energy being particularly significant is also logical, as these properties are closely related to the overall connectivity strength and the compactness of connections within subnetworks. The outcomes for both the ASD and the MDD datasets showed promising results that should be discussed more deeply with neuroscientists. The results mainly suggest the importance of graph substructures or spectral and small-world properties over more basic graph properties to explain how graph neural networks predict these neurological disorders in the FC matrices of patients. This is a result which, to the best of our knowledge, has never been investigated before.